OpenStudy (anonymous):

I really really need help :'( I'm about to cry I'm so frustrated! Algebra 2 help please! 5∑n=1 (3n+1). I need the number of terms, the first, last, and evaluation of the sum :(

OpenStudy (campbell_st):

is the question \[\sum_{n=1}^{5} 3n + 1\]

OpenStudy (anonymous):

I need the first term, the last term, evaluation of the sum and the number of terms :(

OpenStudy (anonymous):

Yes it is

OpenStudy (campbell_st):

ok... so all you do is substitute \[n = 1 .... 3 \times 1 + 1 = 4\] \[n = 2 .... 3\times 2 + 1 = 7\] \[n = 3.... 3 \times 3 + 1 = 10\] \[n = 4 .... 3 \times 4 + 1 = 13\] \[n = 5 .... 3 \times 5 + 1 = 16\] then just add the answers 4 + 7 + 10 + 13 + 16 = hope this helps

OpenStudy (campbell_st):

so if you need 1st term a = 4 last term l = 16 the sum can be found using \[S_{5} = \frac{5}{2} \times( a + l) ..... =.... \frac{5}{2} \times ( 4 + 16)\] this is the more algebraic approach.

OpenStudy (anonymous):

Can you tell me how the first term is 4?

OpenStudy (anonymous):

And also how do you know when you get to the last term, do you just go to 5?

OpenStudy (campbell_st):

ok... so you have an equation that is \[3 \times n + 1 \] substitute n = 1 which is the 1st term in the series... since the values of n start at 1 bottom of the sigma sign and finish at 5... top of the sigma sign. so substituting you get \[3 \times 1 + 1 = 3 + 1 = 4\] hope this helps.

OpenStudy (anonymous):

Because I'm working on my next one, 8∑n=1 (2n/3) and I'm getting my first term as .66666666666667

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

And thank you so much for your help, it was GREATLY appreciated!!!!!!!!!!!

OpenStudy (anonymous):

And my last term came out as 5.33333?

OpenStudy (campbell_st):

ok... so this has 8 terms so to find the terms substitute n = 1 n = 1 2/3 n = 2 4/3 n=3 6/3 . . . . n = 8 16/6

OpenStudy (campbell_st):

oops should be n = 8 16/3 I'd leave all the terms as fractions... its just eaiser to add so the sum = 8/2( 2/3 + 16/3) 4(18/3) Sum = 24

OpenStudy (anonymous):

Thank you so much, I completely understand, really I can not thank you enough :)